# iclr2025_conference.bib

### v3

@@ 4 Proposed SIDE Method
In this section, we will construct implicit informative labels for unconditional DPMs, convert the implicit labels into explicit ones, and then leverage the explicit labels to extract training data.
@@ 4.1 Constructing Implicit Informative Labels
Intuitively, one could use a classifier to generate (predict) implicit labels $y_{I}$ during the sampling process of the diffusion model. The classifier can be a normal classifier trained on the same dataset as the diffusion model. When such a classifier is not available, random labels or representation clusters extracted by a pre-trained feature extractor (e.g., the CLIP image encoder) can also be used as the implicit labels, according to our theoretical analysis in Section 3.2. We assume an implicit label $y_{I}$ is learned by the target unconditional DPM, with its sampling process defined as:
$\mathrm{d}{\bm{x}}=\left[f({\bm{x}},t)-g(t)^{2}\left(\nabla_{\bm{x}}\log p_{\theta}^{t}({\bm{x}})+\nabla_{\bm{x}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\right)\right]\mathrm{d}t+g(t)\mathrm{d}w,$ (10)
@@
We can use the classifier that generates the implicit labels to approximate the gradient in equation (10). However, a known challenge associated with neural network classifiers is their tendency towards miscalibration guo2017calibration. Specifically, the classifier could be overconfident or underconfident about its predictions. To mitigate the potential impact of miscalibration on the sampling process, we introduce a hyperparameter $\lambda$ to calibrate the classifier’s probability output on the diffusion path using power prior as follows:
$p^{t}_{\theta}\left({\bm{x}}|y_{I}\right)\propto p^{t\lambda}_{\theta}\left(y_{I}|{\bm{x}}\right)p^{t}_{\theta}\left({\bm{x}}\right).$ (11)
@@
$\mathrm{d}{\bm{x}}=\left[f({\bm{x}},t)-g(t)^{2}\left(\nabla_{{\bm{x}}}\log p_{\theta}^{t}({\bm{x}})+\lambda\nabla_{{\bm{x}}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\right)\right]\mathrm{d}t+g(t)\mathrm{d}w.$ (12)
@@ 4.2 Time-dependent Classifier
In equation (151), the classifier is denoted by $p_{\theta}^{t}(y|{\bm{x}})$ with $t$ being the timestep, implying its time-dependent nature. However, we do not have a time-dependent classifier but only a time-independent classifier. To address this problem, we propose a method named Time-Dependent Knowledge Distillation (TDKD) to train a time-dependent classifier. The distillation process is illustrated in Figure 2. TDKD equips the classifier with time-dependent guidance during sampling. It operates in two steps: first, the network architecture is adjusted to accommodate time-dependent inputs, and the structure of the time-dependent module and modification are illustrated in Appendix C; second, a synthetic dataset and pseudo labels are created to facilitate knowledge distillation from the normal classifier to its time-dependent counterpart.
As the original training dataset is unknown, we employ the target DPM to generate a synthetic dataset, following the generative data augmentation techniques chen2023data; chen2024comprehensive. We then use the normal classifier trained on the original dataset to generate pseudo labels for the synthetic dataset. Finally, we train a time-dependent classifier on the labelled synthetic dataset. The objective of this training is defined as following:
$\mathcal{L}_{distil}=D_{KL}\left(p_{\theta}(y_{I}|{\bm{x}}),p_{\theta}^{t}(y_{I}|{\bm{x}}_{t})\right).$ (13)
@@
With the trained time-dependent classifier $p_{\theta}^{t}(y|{\bm{x}}_{t})$ and the target DPM, our SIDE extracts training data from the model following a conditional generation process. Assume we condition on the label $y=c$. First, we pick a set of values for $\lambda$, called $\mathcal{S}_{\lambda}$, to use in the SIDE attack. Then, we gather $N_{G}$ data samples for each value of $\lambda$ from the set. During each sampling timestep $t$, we compute the gradient $\nabla_{{\bm{x}}_{t}}CE(c,p_{\theta}^{t}(y|{\bm{x}}_{t}))$ ($CE(\cdot)$ is the cross-entropy loss), then we use the gradient and the target DPM to reverse the diffusion process. Third, we compute the similarity score for each generated image. Lastly, we evaluate the attack performance using evaluation metrics. Defer to Appendix E for SIDE’s pseudocode.
@@ 5 Experiments
In this section, we first describe our experimental setup, introduce the performance metrics, and then present the main evaluation results of our SIDE method. We empirically verify that the memorized images are not from the classifier. We also conduct an ablation study and hyper-parameter analysis to help understand the working mechanism of SIDE.
@@ 5.1 Experimental Setup
We evaluate the effectiveness of our method on three datasets: 1) CelebA-HQ-Face-Identity (CelebA-HQ-FI) na2022unrestricted which consists of 5478 images, 2) a subset of the CelebA (CelebA-25000) liu2015faceattributes which contains 25,000 images and 2) CIFAR-10 containing 50,000 images. All the images are normalized to [-1,1]. We use the AdamW optimizer loshchilov2018decoupled with a learning rate of 1e-4 to train the time-dependent classifier. We train denoising DPMs with a discrete denoising scheduler (DDIM song2021denoising) using the HuggingFace implementation diffusers. All DPMs are trained with batch size 64. We train the models for 258k steps ($\approx$ 3000 epochs) on CelebA-HQ-FI, 390k steps ($\approx$ 1000 epochs) on CelebA-25000 and 1600k steps ($\approx$ 2048 epochs) on CIFAR-10, respectively. We use ResNet34 He2015DeepRL as the normal classifier.
@@ 5.2 Performance Metrics
Determining whether a generated image is a memorized copy of a training image is difficult, as $L_{p}$ distances in the pixel space are ineffective. Previous research addresses this by using the 95th percentile Self-Supervised Descriptor for Image Copy Detection (SSCD) score for image copy detection Somepalli2022DiffusionAO; Gu2023OnMI. However, the 95th percentile SSCD score has three limitations: 1) it does not measure the uniqueness of memorized images; 2) it may underestimate the number of memorized samples when cut at the 95th percentile; and 3) it does not account for different types of memorization. Here, we propose two new memorization scores to solve these issues: 1) Average Memorization Score (AMS) and 2) Unique Memorization Score (UMS metrics).
@@
$AMS\left({{\mathcal{D}}}_{{\text{gen}}},{\mathcal{D}}_{\text{train}},\alpha,\beta\right)=\frac{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{\text{gen}}}{\mathcal{F}\left({\bm{x}}_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta\right)}}{N_{G}},$ (14)
$UMS\left({{\mathcal{D}}}_{{\text{gen}}},{\mathcal{D}}_{\text{train}},\alpha,\beta\right)=\frac{|\bigcup_{{\bm{x}}_{i}\in{\mathcal{D}}_{\text{gen}}}\phi\left({\bm{x}}_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta\right)|}{N_{G}},$ (15)
where ${{\mathcal{D}}}_{\text{gen}}$ is the generated dataset, ${\mathcal{D}}_{\text{train}}$ is the training dataset, and $\alpha$, $\beta$ are thresholds for image similarity scoring. $\mathcal{F}({\bm{x}}_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta)$ serves as a binary check for whether any training sample meets the similarity/distance criteria. $\phi({\bm{x}}_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta)$ provides the specific indices of those training samples that meet the similarity/distance criteria. Mathematically, we can represent $\mathcal{F}(x_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta)=\mathbb{1}[\max_{x_{j}\in{\mathcal{D}}_{\text{train}}}\gamma(x_{i},x_{j})\geq\alpha\;\&\;\gamma(x_{i},x_{j})\leq\beta]$ $\varphi(x_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta)=\{j:x_{j}\in{\mathcal{D}}_{\text{train}},\gamma(x_{i},x_{j})\geq\alpha\;\&\;\gamma(x_{i},x_{j})\leq\beta\}$. $\gamma$ represents the similarity/distance function. For low-resolution datasets, we use the normalized $L_{2}$ distance as $\gamma$ following carlini2023extracting, while for high-resolution datasets, we use the SSCD score as $\gamma$. In our experiments, the thresholds for SSCD are set to $\alpha=0.4$ and $\beta=0.5$ for low similarity, $\alpha=0.5$ and $\beta=0.6$ for mid similarity, and $\alpha=0.6$ and $\beta=1.0$ for high similarity. The thresholds for the normalized $L_{2}$ ditance carlini2023extracting are set to $\alpha=1.5$ and $\beta=10$ for low similarity, $\alpha=1.4$ and $\beta=1.5$ for mid similarity, and $\alpha=1.35$ and $\beta=1.4$ for high similarity.
@@
While carlini2023extracting; chen2024towards introduced metrics similar to AMS and UMS, they did not account for varying levels of similarity, which is essential for assessing different types of copyright infringement, such as character or style copying lee2023talkin; sag2023copyright; sobel2023elements. Additionally, the UMS considers the number of generated images $N_{G}$, which was overlooked in carlini2023extracting. The significance of $N_{G}$ lies in its non-linear impact on the UMS (see Appendix B for proof), indicating that UMS scores should not be compared across different values of $N_{G}$.
@@
5.3 Main Results
We compare SIDE with a random baseline and a variant of SIDE that substitutes the time-dependent classifier with a standard (time-independent) classifier. Here, "TD" refers to the time-dependent classifier trained using our proposed TDKD method, "TI" denotes the time-independent classifier, and "OL" indicates training with the original dataset labels.
The "Random" baseline generates images directly using the target unconditional DPM, as described in carlini2023extracting. We average the results across various values of $\lambda$ (defined in equation (151)), ranging from 5 to 9, with a detailed analysis provided in Section 5.3. It is important to note that $\lambda=0$ corresponds to the "Random" baseline.
For each $\lambda$, including $\lambda=0$, we generate 50,000 images for CelebA and 10,000 images for CIFAR-10 to validate our theoretical analysis and the proposed SIDE method. This effort results in one of the largest generated image datasets to date for studying the memorization of DPMs.
Table 1: The AMS (%) and UMS (%) results at low (top), mid (middle), and high (bottom) levels. ‘Random’ denotes the baseline that generates images directly using the target unconditional DPM, while OL-TI is a variant of SIDE which is not trained using TDKD.
@@ 5.3 Main Results
The AMS and UMS results for the three datasets are presented in Table 1. As shown, SIDE is highly effective in extracting memorized data across all similarity levels, particularly on the CelebA-25000 dataset, a task previously deemed unfeasible due to its large scale.
@@
As shown in Table  1, time-independent classifiers perform significantly worse than their time-dependent counterparts on the two high-resolution CelebA datasets, achieving only about 10% of the effectiveness of classifiers trained using our TDKD method. This discrepancy arises because time-independent classifiers can provide accurate gradients only at the final timestep, whereas time-dependent classifiers deliver accurate gradients at each timestep. However, the performance gap narrows on the CIFAR-10 dataset due to its simplicity, consisting of only 10 classes. Thus, the gradients produced by the time-independent classifier are less prone to inaccuracy, which mitigates the limitations of the time-independent classifier.
@@
There might be a concern that the extracted images may originate from the classifier rather than the target DPM. We argue that even if a time-dependent classifier could disclose its training images, these images would still originate from the target DPM, as they can also be extracted with the time-dependent classifiers. Furthermore, Table  1 demonstrates that employing a simple classifier reduces memorization compared to the baseline extraction method, indicating that the memorized images do not originate from the classifier.
@@
The classifier used to train SIDE is associated with a certain number of classes. Here, we conduct experiments to explore the relationship between the number of classes (of the classifier) and the extraction performance at a low similarity level, using 1,200 images per class with $\lambda=5$ on the CelebA-HQ-FI dataset. As shown in Table  2 and Figure 7 (Appendix), there exists a strong positive correlation: as the number of classes increases, both AMS and UMS improve. In summary, increasing number of classes positively affects both AMS and UMS, with a stronger impact on AMS. The UMS values reported here differ significantly from Table  1. This is because here, we only generated 1,200 images per class, whereas 50,000 images per class in the previous experiment. Increasing the number of classes improves AMS more than UMS because a higher class count enables the classifier to better differentiate fine details, leading to more accurate matches at low similarity levels. UMS is less affected since it relies more on the diversity across images, which is constrained by the smaller number of images per class in this experiment.
Class and AMS 7.4 0.115 0.637 0.80
Class and UMS 6.1 0.090 0.483 0.70
@@
Here, we test the sensitivity of SIDE to its hyper-parameter $\lambda$. We generate 50,000 images for each integer value of $\lambda$ within the range of [0, 50]. As shown in Figure 4, the memorization score increases at first, reaching its highest, then decreases as $\lambda$ increases. This can be understood from sampling SDE Equation (151). Starting from 0, the diffusion models are unconditional. As $\lambda$ increases, the diffusion models become conditional, and according to Theorem 2, the memorization effect will be triggered. However, when $\lambda$ is too large, the generated images will overfit the classifier’s decision boundaries, leading to a low diversity and ignoring the data distribution. Consequently, the memorization score decreases.

### v5

@@ 3 Proposed SIDE Method
3.1 Intuition
Previous studies [13, 32] have shown that conditional DPMs are more prone to memorization due to their reliance on explicit conditioning labels, such as class labels or text conditions, which we refer to as explicit informative labels or explicit labels. In contrast, unconditional DPMs do not have access to these explicit labels. Nevertheless, they still learn to form representation clusters from the training data, even though these clusters are not directly specified in the data itself [8]. We name these representation clusters as implicit informative labels or implicit labels. By formalizing the clustering patterns within the training data, it may be possible to construct implicit labels and make the unconditional DPMs conditional on these implicit labels to extract training data from unconditional DPMs.
In the following section, we will construct implicit informative labels for unconditional DPMs, convert the implicit labels into explicit ones, and then leverage the explicit labels to extract training data.
@@ 3.2 Constructing Implicit Informative Labels
Intuitively, one could use a classifier to generate (predict) implicit labels $y_{I}$ during the sampling process of the diffusion model. The classifier can be a normal classifier trained on the same dataset as the diffusion model. When such a classifier is not available, random labels or representation clusters extracted by a pre-trained feature extractor (e.g., the CLIP image encoder) can also be used as the implicit labels. We assume an implicit label $y_{I}$ is learned by the target unconditional DPM, with its sampling process:
$\mathrm{d}{\bm{x}}=$ $\Big{[}f({\bm{x}},t)-g(t)^{2}\Big{(}\nabla_{\bm{x}}\log p_{\theta}^{t}({\bm{x}})+$
$\nabla_{\bm{x}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\Big{)}\Big{]}\mathrm{d}t+g(t)\mathrm{d}w$ (1)
@@
We can use the classifier that generates the implicit labels to approximate the gradient in Eq. (3.2). However, a known challenge associated with neural network classifiers is their tendency towards miscalibration [14]. Specifically, the classifier could be overconfident or underconfident about its predictions. To mitigate the potential impact of miscalibration on the sampling process, we introduce a hyperparameter $\lambda$ to calibrate the classifier’s probability output on the diffusion path using power prior as follows:
$p^{t}_{\theta}\left({\bm{x}}|y_{I}\right)\propto p^{t\lambda}_{\theta}\left(y_{I}|{\bm{x}}\right)p^{t}_{\theta}\left({\bm{x}}\right)$ (2)
@@
$\mathrm{d}{\bm{x}}=$ $\Big{[}f({\bm{x}},t)-g(t)^{2}\big{(}\nabla_{{\bm{x}}}\log p_{\theta}^{t}({\bm{x}})$
$+\lambda\nabla_{{\bm{x}}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\big{)}\Big{]}\mathrm{d}t+g(t)\mathrm{d}w$ (3)
@@ 3.3 Time-dependent Classifier
In Eq. (3.2), the classifier is denoted by $p_{\theta}^{t}(y|{\bm{x}})$ with $t$ being the timestep, implying its time-dependent nature. However, we do not have a time-dependent classifier but only a time-independent classifier. To address this problem, we propose a method named Time-Dependent Knowledge Distillation (TDKD) to train a time-dependent classifier. The distillation process is illustrated in Figure 2. TDKD equips the classifier with time-dependent guidance during sampling. It operates in two steps: first, the network architecture is adjusted to accommodate time-dependent inputs, and the structure of the time-dependent module and modification are illustrated in Appendix C; second, a synthetic dataset and pseudo labels are created to distil knowledge from the normal classifier to its time-dependent counterpart.
As the original training dataset is unknown, we employ the target DPM to generate a synthetic dataset, following the generative data augmentation techniques [10, 9]. We then use the normal classifier trained on the original dataset to generate pseudo labels for the synthetic dataset. Finally, we train a time-dependent classifier on the labelled synthetic dataset. The objective of this training is defined as follows:
$\mathcal{L}_{distill}=D_{KL}\left(p_{\theta}(y_{I}|{\bm{x}}),p_{\theta}^{t}(y_{I}|{\bm{x}}_{t})\right)$ (4)
@@
With the trained time-dependent classifier $p_{\theta}^{t}(y|{\bm{x}}_{t})$ and the target DPM, our SIDE extracts training data from the model following a conditional generation process. Assume we condition on the label $y=c$. Firstly, we pick a set of values for $\lambda$, called $\mathcal{S}_{\lambda}$, to use in the SIDE attack. Then, we gather $N_{G}$ data samples for each value of $\lambda$ from the set. For each sampling timestep $t$, we compute the gradient $\nabla_{{\bm{x}}_{t}}CE(c,p_{\theta}^{t}(y|{\bm{x}}_{t}))$ ($CE(\cdot)$ denotes the cross-entropy loss) and use the gradient and the target DPM to reverse the diffusion process. Next, we compute the similarity score for each generated image. Finally, we evaluate the attack performance using evaluation metrics. Defer to Appendix E for SIDE’s pseudocode.
@@ 4 Experiments
In this section, we first describe our experimental setup, introduce the performance metrics and then present the main evaluation results of our SIDE method. We empirically verify that the memorized images are not from the classifier. We also conduct an ablation study and hyper-parameter analysis to help understand the working mechanism of SIDE.
@@ 4.1 Experimental Setup
We evaluate the effectiveness of our method on three datasets: 1) CelebA-HQ-Face-Identity (CelebA-HQ-FI) [23] which consists of 5478 images, 2) a subset of the CelebA (CelebA-25000) [21] which contains 25,000 images and 2) CIFAR-10 containing 50,000 images. All the images are normalized to [-1,1]. We use the AdamW optimizer [22] with a learning rate of 1e-4 to train the time-dependent classifier. We train denoising DPMs with a discrete denoising scheduler (DDIM [33]) using the HuggingFace implementation [36]. All DPMs are trained with batch size 64. We train the models for 258k steps ($\approx$ 3000 epochs) on CelebA-HQ-FI, 390k steps ($\approx$ 1000 epochs) on CelebA-25000 and 1600k steps ($\approx$ 2048 epochs) on CIFAR-10, respectively. We use ResNet34 [15] as the normal classifier.
@@ 4.2 Performance Metrics
Determining whether a generated image is a memorized copy of a training image is difficult, as $L_{p}$ distances in the pixel space are ineffective. Previous research addresses this by using the 95th percentile Self-Supervised Descriptor for Image Copy Detection (SSCD) score for image copy detection [31, 13]. However, the 95th percentile SSCD score has three limitations: 1) it does not measure the uniqueness of memorized images; 2) it may underestimate the number of memorized samples when cut at the 95th percentile; and 3) it does not account for different types of memorization. Here, we propose two new memorization scores to solve these issues: 1) Average Memorization Score (AMS) and 2) Unique Memorization Score (UMS).
@@
$AMS\left({{\mathcal{D}}}_{1},{\mathcal{D}}_{2},\alpha,\beta\right)=\frac{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{1}}{\mathcal{F}\left({\bm{x}}_{i},{\mathcal{D}}_{1},\alpha,\beta\right)}}{N_{G}}$ (5)
$UMS\left({{\mathcal{D}}}_{1},{\mathcal{D}}_{2},\alpha,\beta\right)=\frac{|\bigcup_{{\bm{x}}_{i}\in{\mathcal{D}}_{1}}\phi\left({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta\right)|}{N_{G}}$ (6)
where ${{\mathcal{D}}}_{1}$ is the generated dataset, ${\mathcal{D}}_{2}$ is the training dataset, and $\alpha$, $\beta$ are thresholds for image similarity scoring. $\mathcal{F}({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta)$ serves as a binary check for whether any training sample meets the similarity/distance criteria. $\phi({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta)$ provides the specific indices of those training samples that meet the similarity/distance criteria. Mathematically, we can represent:
$\mathcal{F}(x_{i},{\mathcal{D}}_{2},\alpha,\beta)=\mathds{1}\Big{[}\max_{x_{j}\in{\mathcal{D}}_{2}}\gamma(x_{i},x_{j})\in[\alpha,\beta]\Big{]}$ (7)
$\phi(x_{i},{\mathcal{D}}_{2},\alpha,\beta)=\{j:x_{j}\in{\mathcal{D}}_{2},\gamma(x_{i},x_{j})\in[\alpha,\beta]\}$ (8)
where $\gamma$ represents the similarity/distance function. For low-resolution datasets, we use the normalized $L_{2}$ distance as $\gamma$ following [5], while for high-resolution datasets, we use the SSCD score as $\gamma$. In our experiments, the thresholds for SSCD are set to $\alpha=0.4$ and $\beta=0.5$ for low similarity, $\alpha=0.5$ and $\beta=0.6$ for mid similarity, and $\alpha=0.6$ and $\beta=1.0$ for high similarity. (examples of different thresholds are shown in Figure 3) The thresholds for the normalized $L_{2}$ ditance [5] are set to $\alpha=1.5$ and $\beta=10$ for low similarity, $\alpha=1.4$ and $\beta=1.5$ for mid similarity, and $\alpha=1.35$ and $\beta=1.4$ for high similarity.
@@
While [5, 7] introduced metrics similar to AMS and UMS, they did not account for varying levels of similarity, which is essential for assessing different types of copyright infringement, such as character or style copying [20, 27, 29]. Additionally, the UMS considers the number of generated images $N_{G}$, which was overlooked in [5]. The significance of $N_{G}$ lies in its non-linear impact on the UMS (see Appendix B for proof), indicating that UMS scores should not be compared across different values of $N_{G}$.
@@
Table 1: The AMS (%) and UMS (%) results at low, mid, and high levels. ‘Random’ denotes the baseline that generates images directly using the target unconditional DPM[5], while OL-TI is a variant of SIDE which is not trained using TDKD.
@@
4.3 Main Results
We compare SIDE with the Random baseline proposed by Carlini et al.[5] and a variant of SIDE that substitutes the time-dependent classifier with a standard (time-independent) classifier. Here, ”TD” refers to the time-dependent classifier trained using our proposed TDKD method, ”TI” denotes the time-independent classifier, and ”OL” indicates training with original dataset labels.
The ”Random” baseline generates images directly using the target unconditional DPM, as described in [5]. We average the results across various values of $\lambda$ (defined in Eq. (3.2)), ranging from 5 to 9, with a detailed analysis provided in Section 4.3. It is important to note that $\lambda=0$ corresponds to the ”Random” baseline.
For each $\lambda$, including $\lambda=0$, we generate 50,000 images for CelebA and 10,000 images for CIFAR-10 to validate the proposed SIDE method. This effort results in one of the most enormous generated image datasets for studying the memorization of DPMs.
#Class vs. AMS 7.4 (positive) 0.115 0.637 0.80
#Class and UMS 6.1 (positive) 0.090 0.483 0.70
@@ 4.3 Main Results
The AMS and UMS results for the three datasets are presented in Table 1 (results of 95-th percentile of SSCD are listed in Appendix F). As shown, SIDE is highly effective in extracting memorized data across all similarity levels, particularly on the CelebA-25000 dataset, a task previously deemed unfeasible due to its large scale.
@@
As shown in Table 1, time-independent classifiers perform significantly worse than their time-dependent counterparts on the two high-resolution CelebA datasets, achieving only about 10% of the effectiveness of classifiers trained using our TDKD method. This discrepancy arises because time-independent classifiers can provide accurate gradients only at the final timestep, whereas time-dependent classifiers deliver accurate gradients at each timestep. However, the performance gap narrows on the CIFAR-10 dataset due to its simplicity, consisting of only 10 classes. Thus, the gradients produced by the time-independent classifier are less prone to inaccuracy, which mitigates the limitations of the time-independent classifier.
@@
There might be a concern that the extracted images may originate from the classifier rather than the target DPM. We argue that even if a time-dependent classifier could disclose its training images, these images would still originate from the target DPM because the classifier is distilled on the synthetic dataset generated by the target DPM. Furthermore, Table  1 demonstrates that employing a simple classifier reduces memorization compared to the baseline extraction method, indicating that the memorized images do not originate from the simple classifier as well.
@@
The classifier used to train SIDE is associated with a certain number of classes. Here, we conduct experiments to explore the relationship between the number of classes (of the classifier) and the extraction performance at a low similarity level, using 1,200 images per class with $\lambda=5$ on the CelebA-HQ-FI dataset. As shown in Table  2 and Figure 5, there exists a strong positive correlation: as the number of classes increases, both AMS and UMS improve. In summary, increasing number of classes positively affects both AMS and UMS, with a stronger impact on AMS. The UMS values reported here differ significantly from Table  1. This is because we only generated 1,200 images per class here, whereas 50,000 images per class in the previous experiment. Increasing the number of classes improves AMS more than UMS because a higher class count enables the classifier to better differentiate fine details, leading to more accurate matches at low similarity levels. UMS is less affected since it relies more on the diversity across images, which is constrained by the smaller number of images per class in this experiment.
@@
Here, we test the sensitivity of SIDE to its hyper-parameter $\lambda$. We generate 50,000 images for each integer value of $\lambda$ within the range of [0, 50]. As shown in Figure 4, the memorization score increases at first, reaching its highest, then decreases as $\lambda$ increases. This can be understood from sampling SDE Eq. (3.2). Starting from 0, the diffusion models are unconditional. As $\lambda$ increases, the diffusion models become conditional, and according to Theorem 1, the memorization effect will be triggered. However, when $\lambda$ is too large, the generated images will overfit the classifier’s decision boundaries, leading to a low diversity and ignoring the data distribution. Consequently, the memorization score decreases.
5 Theoretical Understanding for Memorization in Conditional Generative Models
In this section, we introduce a metric to quantify the memorization capacity of generative models and provide a theoretical explanation for the data memorization in conditional diffusion models and why SIDE is effective.
5.1 Memorization Metric
Intuitively, the memorization of fixed training data points (i.e., point-wise memorization) can be quantified by the degree of overlap between the generated distribution and the distributions centered at each training data point. Given a generative model and training dataset, we propose the following memorization metric to quantify the memorization of the training data points in the model.
Definition 1 (Point-wise Memorization).
Given a generative model $p_{\theta}$ with parameters $\theta$ and training dataset ${\mathcal{D}}=\{{\bm{x}}_{i}\}_{i=1}^{N}$, degree of memorization in $p_{\theta}$ of ${\mathcal{D}}$ is defined as:
${\mathcal{M}}_{p}({\mathcal{D}};p_{\theta})=\frac{1}{|{\mathcal{D}}|}\sum_{{\bm{x}}_{i}\in{\mathcal{D}}}{\int{p_{\theta}}}({\bm{x}})\log\frac{p_{\theta}({\bm{x}})}{q({\bm{x}},{\bm{x}}_{i},\epsilon)}\,d{\bm{x}}$ (9)
where ${\bm{x}}_{i}\in\mathbb{R}^{d}$ denotes the $i$-th training sample, $N$ is the total number of training samples, $p_{\theta}({\bm{x}})$ represents the probability density function (PDF) of the generated samples, and $q({\bm{x}},{\bm{x}}_{i},\epsilon)$ is the probability distribution centered at training sample ${\bm{x}}_{i}$ within a small radius $\epsilon$.
A straightforward choice for $q({\bm{x}},{\bm{x}}_{i},\epsilon)$ is the Dirac delta function centered at training data point ${\bm{x}}_{i}$: $q({\bm{x}},{\bm{x}}_{i},\epsilon)=\delta({\bm{x}}-{\bm{x}}_{i})$. However, this would make Eq. (9) uncomputable as the Dirac delta function is zero beyond the $\epsilon$-neighborhood. Alternatively, we could let $q({\bm{x}},{\bm{x}}_{i},\epsilon)\sim{\mathcal{N}}(x_{i};\epsilon I)$. Note that in Eq. (9), a smaller value of ${\mathcal{M}}_{p}({\mathcal{D}};\theta)$ indicates more overlap between the two distributions and thus more memorization. As $\epsilon$ tends to 00, the measured memorization becomes more accurate, with the limit characterizing the memorization capability of $p_{\theta}$.
5.2 Theoretical Analysis
Based on the point-wise memorization metric, we perform a theoretical understanding that explains why conditional generative models achieve better memorization capacity. Our analysis is based on the concept of informative labels, which refers to information that can partition a dataset into various disjoint subsets. And we demonstrate that provided with an informative label, the conditional generative models achieve stronger memorization capacity.
Informative Labels
The concept of informative labels has been previously discussed as class labels [13]. In this work, we introduce a more general definition that encompasses both class labels and random labels as special cases. We formally define an informative label as follows.
Definition 2 (Informative Label).
Let ${\mathcal{Y}}$ be an attribute of the data taking values in $\{y_{i}\}_{i=1}^{C}$. We refer to ${\mathcal{Y}}$ as an informative label if we can partition the dataset into several mutually disjoint subsets $\{{\mathcal{D}}_{i}\}_{i=1}^{C}$ based on the values of the attribute ${\mathcal{Y}}$.
Here, the informative labels are not limited to the conventional class labels; they can also be text captions, features, or cluster information that can be used to group the training samples into subsets. The above definition states that an information label should be able to differentiate a subset of samples from others. An extreme case is that all samples have the same label; in this case, the label is not informative. According to our definition, class-wise and random labels are special cases of informative labels. Informative labels can be either explicit like class/random labels and text captions, or implicit like salient features or clusters.
Main Results
Here we introduce our main theorem on the memorization mechanism of conditional generative models and explain why SIDE is effective. We use ${\mathcal{D}}_{i}$ to denote the subset consisting of data with informative label ${\mathcal{Y}}=y_{i}$. In addition, we denote the distribution of data in the original dataset ${\mathcal{D}}$ by $p$. The distribution has expectation $\mu$ and covariance matrix $\Sigma$. Similarly, for each attribute $y_{i}$, we use $p_{i}$ to denote the distribution of the data with this attribute, and use $\mu_{i}$ and $\Sigma_{i}$ to denote its expectation and covariance matrix.
Theorem 1.
If a generative model $p_{\theta_{i}}$ matches the target distribution $p_{i}$ well for each $i$, i.e. $TV(p_{i},p_{\theta_{i}})=0$, and the data distribution for some informative label $y_{c}$ satisfies $tr(\Sigma_{c})\leq||\mu-\mu_{c}||_{2}^{2}$, we have with probability 1:
$\lim_{|{\mathcal{D}}_{c}|\to\infty}\lim_{\epsilon\to 0}\frac{{\mathcal{M}}_{p}({\mathcal{D}}_{c},p_{\theta_{c}})}{{\mathcal{M}}_{p}({\mathcal{D}}_{c},p_{\theta})}\leq 1$ (10)
where $p_{\theta_{i}}$ denotes the distribution of generated data conditioned on ${\mathcal{Y}}=y_{i}$.
The proof can be found in Appendix A. Theorem 1 states that any form of information labels can enhance memorization capacity in conditional generative models, including conventional class labels and random labels. The condition $tr(\Sigma_{c})\leq||\mu-\mu_{c}||_{2}^{2}$ is reasonable because it ensures that the spread (variance) of the data within each class is not excessively large.
$\frac{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{c}}{\left({\bm{x}}_{i}-\mu_{c}\right)^{T}\left({\bm{x}}_{i}-\mu_{c}\right)}}{|{\mathcal{D}}_{c}|}\leq||\mu-\mu_{c}||_{2}^{2}$ (11)
To understand this, we look at the derivation Eq. (11) from $tr(\Sigma_{c})\leq||\mu-\mu_{c}||_{2}^{2}$. For a well-separated class, this squared distance should provide an upper bound on how spread out the class data can be. If $tr(\Sigma_{c})$ were large relative to $||\mu-\mu_{c}||_{2}^{2}$, it would imply that data samples represented by informative label $c$ are highly spread out, making it harder for the class labelled with $c$ to be well-formed and distinct. Thus, the condition ensures that the data samples form tight clusters around their means, making it reasonable for the class to be well-defined and easily detachable.
This also explains why SIDE works. Chen et al. [8] found that unconditional DPMs learn clustered representations from training data. The data samples within each class should not be widely dispersed for effective clustering. Then, the samples should satisfy Eq. 11. So, although these clustered representations are implicit, we could extract training data from unconditional DPMs if we can convert them from implicit informative labels into explicit labels. When conditioning on the explicit labels, the DPMs will exhibit more memorization effects indicated by Theorem 1, which makes room for data extraction attacks.
6 Limitations:
SIDE has certain limitations that warrant further research. SIDE mainly uses human-labelled data to extract data, which is scarce and limits its practical use. Future work could investigate hybrid approaches for generating implicit labels, such as employing unsupervised clustering or utilizing pretrained models to label synthetic datasets.